Calculate the overall percent yield of the synthesis to produce crude lidocaine.

3.0ml of 2,6-dimethylaniline and 2.0ml of alpha-chloroacetyl was used to produce 8.349g of alpha chloro 2,6 dimethylacetanilide. 7.5 ml diethylamine was added to the alpha chloro 2,6 dimethylacetanilide and 2.391g of crude lidocaine was produced.
Answer #1

Step-1:

2,6-dimethylaniline density = 0.9842 g/mL

mass = 0.9842 g/mL x 3 mL = 2.9526 g

moles = 2.9526 g / 121.18 g/mol = 0.02436540683 mol

alpha-chloroacetyl density = 1.4202 g/mL

mass = 1.4202 g/mL x 2 mL = 2.8404 g

moles = 2.8404 g / 112.94 g/mol = 0.02514963697 mol

Theoretical moles of alpha chloro 2,6 dimethylacetanilide = 0.02436540683 mol

Theoretical yield = 0.02436540683 mol x 197.66 g/mol = 4.81606631402 g

Percent yield = ( 8.349 g / 4.81606631402 ) x 100% = 173.36 % ----- This is impossible, but i am calculated, percent yield never cross 100% in any reaction

------------------------------------------------------------------------------------------------

Step-2:

moles of alpha chloro 2,6 dimethylacetanilide = 0.02436540683 mol

diethylamine density = 0.7056 g/mL

mass = 0.7056 g/mL x 7.5 mL = 5.292 g

moles = 5.292 g / 73.14 g/mol = 0.07235438884 mol

but 0.02436540683 moles only reacts

moles lidocaine formed = 0.02436540683 mol

theoretical yield = 0.02436540683 mol x 234.34 g/mol = 5.70978943654 g

percent yield = (2.391 g / 5.70978943654 g) x 100% = 41.87%

------------------------------------------------------------------------------------------------

Overall yield = (173.36 % * 41.87%) = 72.585832 %

Answer: Overall yield = 72.6 %

------------------------------------------------------------------------------------------------

Hope this helped you!

Thank You So Much!

Please Rate this answer as you wish.("Thumbs Up")